### v2

# On a vanishing theorem for surfaces

Abstract.
We propose a new formulation of a vanishing theorem for surfaces. Although this vanishing theorem follows easily from the well-known Kawamata–Viehweg vanishing theorem, it turns out to be remarkably useful. In particular, it is sufficient for the minimal model theory of log surfaces, and it allows one to carry out both the minimal model program and the abundance theorem for log surfaces without invoking any of the deeper vanishing theorems.
Key words and phrases:
vanishing theorems, algebraic surfaces, complex analytic surfaces, minimal model program
2020 Mathematics Subject Classification:
Primary 32C15; Secondary 14E30
1. Introduction
This short paper aims to make the minimal model theory for log surfaces, as established in [F2] and [M], more accessible. In [F2], we develop the minimal model theory for algebraic log surfaces in characteristic zero, building on the framework from [F1]. In contrast, [M] addresses complex analytic log surfaces within the framework of [F5]. Both [F2] and [M] rely on intricate vanishing theorems derived from [F1] and [F6]. In this paper, we demonstrate that the vanishing theorem presented in Theorem 1.1 follows easily from the well-known Kawamata–Viehweg vanishing theorem. This result shows that the Kawamata–Viehweg vanishing theorem is sufficient for the minimal model theory of log surfaces discussed in [F2] and [M]. Consequently, readers primarily interested in the surface theory will be able to avoid the more complicated vanishing theorems of [F1], [F5], and [F6].
Theorem 1.1 (Main Theorem).
Let $X$ be a smooth complex analytic surface and let $\Delta$ be a boundary $\mathbb{R}$-divisor on $X$ such that $\operatorname{Supp}\Delta$ is a simple normal crossing divisor. Let $f\colon X\to Y$ be a bimeromorphic morphism to a complex analytic space $Y$, and let $\pi\colon Y\to Z$ be a proper surjective morphism of complex analytic spaces such that $\psi:=\pi\circ f$ is projective. Let $\mathcal{L}$ be a line bundle on $X$ such that $\mathcal{L}-(\omega_{X}+\Delta)$ is $\psi$-nef and $\psi$-big over $Z$. Assume that $\bigl(\mathcal{L}-(\omega_{X}+\Delta)\bigr)\cdot C>0$ for every irreducible component $C$ of $\lfloor\Delta\rfloor$ such that $f(C)$ is a curve and $\psi(C)$ is a point. Then
$R^{p}\pi_{*}R^{q}f_{*}\mathcal{L}=0$
for every $p>0$ and every $q$.
Theorem 1.1 is not included in the vanishing theorems stated in [F1], [F3, Chapter 5], [F5], or [F6], and, to the best of our knowledge, it appears to be a new formulation. By Theorem 1.1, we can formulate and prove the Kodaira vanishing theorem, the vanishing theorem of Reid–Fukuda type, and related results for log canonical surfaces in both the algebraic and complex analytic settings, without resorting to the intricate vanishing theorems in [F3, Chapter 5], [F6], or [FF]. We leave the details to the interested reader; see [F3, Theorems 5.6.4 and 5.7.6] for further discussion.
The vanishing theorem in [F1] is proved using the theory of mixed Hodge structures on cohomology with compact support. The work in [F5] relies on Saito’s theory of mixed Hodge modules, while [FF] provides an alternative approach to the results of [F5]. As indicated by its title, [FF] employs the theory of variations of mixed Hodge structures.
In contrast, in this paper we only make use of the (relative) Kawamata–Viehweg vanishing theorem. Although its validity for projective morphisms between complex analytic spaces is not entirely obvious, it is widely accepted as a standard result by experts. For readers interested exclusively in algebraic varieties, everything can be assumed algebraic, and the classical Kawamata–Viehweg vanishing can be applied throughout this paper.
We note that the formulation of Theorem 1.1 is new, even in the purely algebraic setting.
Before proceeding, let us briefly outline the structure of this short paper. Section 2 collects the preliminaries: we recall several basic definitions and state the Kawamata–Viehweg vanishing theorem for projective morphisms between complex analytic spaces without proof. In Section 3, we prove two results that follow rather easily from the Kawamata–Viehweg vanishing theorem. Section 4 is devoted to the proof of Theorem 1.1. Finally, in Section 5, we explain, for each relevant paper, that the minimal model theory for log surfaces can be carried out using only the Kawamata–Viehweg vanishing theorem in light of the results established here.
2. Preliminaries
In this paper, we freely use the standard notation and definitions of the minimal model theory for algebraic varieties from [F1], and for projective morphisms of complex analytic spaces from [F5].
Definition 2.1 (Log surfaces).
Let $X$ be a normal complex analytic surface, and let $\Delta$ be a boundary $\mathbb{R}$-divisor on $X$ such that $K_{X}+\Delta$ is $\mathbb{R}$-Cartier. Then the pair $(X,\Delta)$ is called a log surface. We recall that an $\mathbb{R}$-divisor is said to be a boundary $\mathbb{R}$-divisor if all its coefficients lie in the interval $[0,1]$.
We discuss in [F2] and [M] that the minimal model program and the abundance conjecture hold for $\mathbb{Q}$-factorial log surfaces $(X,\Delta)$ without assuming that they are necessarily log canonical.
Definition 2.2 (Nefness).
Let $f\colon X\to Y$ be a projective morphism between complex analytic spaces, and let $\mathcal{L}$ be an $\mathbb{R}$-line bundle, or the sum of an $\mathbb{R}$-line bundle and an $\mathbb{R}$-Cartier divisor, on $X$. We say that $\mathcal{L}$ is $f$-nef over $Y$ if
$\mathcal{L}\cdot C\geq 0$
for every curve $C\subset X$ such that $f(C)$ is a point.
There are various characterizations of relative bigness. The following one is taken from [N2, Chapter II, 5.17. Corollary].
Definition 2.3 (Bigness).
Let $f\colon X\to Y$ be a projective surjective morphism from a normal complex analytic space $X$. Let $\mathcal{L}$ be an $\mathbb{R}$-line bundle, or the sum of an $\mathbb{R}$-line bundle and an $\mathbb{R}$-Cartier divisor, on $X$. We say that $\mathcal{L}$ is $f$-big over $Y$ if, for any irreducible component $F$ of an analytically sufficiently general fiber of $f$, the restriction $\mathcal{L}|_{F}$ is big in the usual sense.
In this paper, we will freely use the following relative Kawamata–Viehweg vanishing theorem for projective morphisms between complex analytic spaces.
Theorem 2.4 (Relative Kawamata–Viehweg vanishing theorem).
Let $f\colon X\to Y$ be a projective surjective morphism from a smooth complex analytic space $X$ and let $\Delta$ be an $\mathbb{R}$-divisor on $X$ such that $\lfloor\Delta\rfloor=0$ and $\operatorname{Supp}\Delta$ is a simple normal crossing divisor. Let $\mathcal{L}$ be a line bundle on $X$ such that $\mathcal{L}-(\omega_{X}+\Delta)$ is $f$-nef and $f$-big over $Y$. Then $R^{i}f_{*}\mathcal{L}=0$ holds for every $i>0$.
For the details of Theorem 2.4, see, for example, [N1, Theorem 3.7] and [N2, Chapter II. 5.12. Corollary].
3. Lemmas
We begin with a vanishing theorem of Reid–Fukuda type for surfaces, which is a slight generalization of the Kawamata–Viehweg vanishing theorem.
Lemma 3.1 (Vanishing theorem of Reid–Fukuda type).
Let $V$ be a smooth complex analytic surface, and let $\Delta$ be a boundary $\mathbb{R}$-divisor on $V$ such that $\operatorname{Supp}\Delta$ is a simple normal crossing divisor. Let $\varphi\colon V\to W$ be a projective surjective morphism of complex analytic spaces, and let $\mathcal{L}$ be a line bundle on $V$ such that $\mathcal{L}-(\omega_{V}+\Delta)$ is $\varphi$-nef and $\varphi$-big over $W$. Assume further that
$\bigl(\mathcal{L}-(\omega_{V}+\Delta)\bigr)\cdot C>0$
for every irreducible component $C$ of $\lfloor\Delta\rfloor$ that is mapped to a point by $\varphi$. Then $R^{i}\varphi_{*}\mathcal{L}=0$ for every $i>0$.
Lemma 3.1 follows easily from the relative Kawamata–Viehweg vanishing theorem (Theorem 2.4), but we provide a proof here for completeness.
Proof of Lemma 3.1.
If $\lfloor\Delta\rfloor=0$, then $R^{i}\varphi_{*}\mathcal{L}=0$ for every $i>0$ by the relative Kawamata–Viehweg vanishing theorem. Thus we may assume that $\lfloor\Delta\rfloor\neq 0$. Without loss of generality, we may further assume that $\lfloor\Delta\rfloor$ has only finitely many irreducible components, since the statement is local over $W$.
Take an irreducible component $C$ of $\lfloor\Delta\rfloor$. Consider the short exact sequence
$0\to\mathcal{L}(-C)\to\mathcal{L}\to\mathcal{L}|_{C}\to 0,$
where $\mathcal{L}(-C):=\mathcal{L}\otimes\mathcal{O}_{V}(-C)$. Note that
$\mathcal{L}(-C)-(\omega_{V}+\Delta-C)=\mathcal{L}-(\omega_{V}+\Delta).$
By induction on the number of irreducible components of $\lfloor\Delta\rfloor$, we obtain $R^{i}\varphi_{*}\mathcal{L}(-C)=0$ for every $i>0$.
Next, observe that
$\mathcal{L}|_{C}-\bigl(\omega_{C}+(\Delta-C)|_{C}\bigr)=\bigl(\mathcal{L}-(\omega_{V}+\Delta)\bigr)|_{C}.$
Since $C$ is a curve, it follows immediately that $R^{i}\varphi_{*}(\mathcal{L}|_{C})=0$ for every $i>0$.
Therefore, the long exact sequence associated to
$0\to\mathcal{L}(-C)\to\mathcal{L}\to\mathcal{L}|_{C}\to 0$
yields $R^{i}\varphi_{*}\mathcal{L}=0$ for all $i>0$. This completes the proof of Lemma 3.1. ∎
The following lemma is a straightforward application of Lemma 3.1. Alternatively, one may use a more elementary argument, as in [F4, Theorem 11.1], to prove Lemma 3.2. In any case, Lemma 3.2 is easy to prove, but it is nevertheless quite useful.
Lemma 3.2.
Let $f\colon X\to Y$ be a projective bimeromorphic morphism from a smooth complex analytic surface $X$, and let $\Delta$ be a boundary $\mathbb{R}$-divisor on $X$ such that $\operatorname{Supp}\Delta$ is a simple normal crossing divisor. Assume that $\mathcal{L}-(\omega_{X}+\Delta)$ is $f$-nef over $Y$.
If $R^{1}f_{*}\mathcal{L}\neq 0$, then the support of $R^{1}f_{*}\mathcal{L}$ is contained in $f(B)$, where $B$ is the union of the irreducible components of $\lfloor\Delta\rfloor$ that are mapped to a point by $f$. In particular, $R^{1}f_{*}\mathcal{L}$ is a skyscraper sheaf on $Y$.
Moreover, $f_{*}\mathcal{L}$ is a torsion-free coherent sheaf on $Y$, and $R^{i}f_{*}\mathcal{L}=0$ for every $i\geq 2$.
Proof of Lemma 3.2.
Consider the short exact sequence
$0\to\mathcal{L}(-B)\to\mathcal{L}\to\mathcal{L}|_{B}\to 0,$
where $\mathcal{L}(-B):=\mathcal{L}\otimes\mathcal{O}_{X}(-B)$. Note that
$\mathcal{L}(-B)-(\omega_{X}+\Delta-B)=\mathcal{L}-(\omega_{X}+\Delta).$
Since $f$ is projective bimeromorphic, $\mathcal{L}-(\omega_{X}+\Delta)$ is $f$-nef and $f$-big over $Y$. By definition, $C\to f(C)$ is finite for any irreducible component $C$ of $\lfloor\Delta-B\rfloor$. Therefore, Lemma 3.1 implies
$R^{i}f_{*}\mathcal{L}(-B)=0\quad\text{for every }i>0.$
The associated long exact sequence then gives
$R^{i}f_{*}\mathcal{L}\simeq R^{i}f_{*}(\mathcal{L}|_{B})=H^{i}(B,\mathcal{L}|_{B})$
for every $i>0$. Hence, the support of $R^{1}f_{*}\mathcal{L}$ is contained in $f(B)$, as claimed. This completes the proof of Lemma 3.2. ∎
4. Proof of Theorem 1.1
In this section, we prove Theorem 1.1. Theorem 1.1 follows directly from Lemmas 3.1 and 3.2, and is thus a slight generalization of the relative Kawamata–Viehweg vanishing theorem (see Theorem 2.4).
Proof of Theorem 1.1.
By Lemma 3.2, it suffices to show that $R^{p}\pi_{*}(f_{*}\mathcal{L})=0$ for every $p>0$.
Let $B$ be the union of the irreducible components of $\lfloor\Delta\rfloor$ that are mapped to a point by $f$, and consider
$0\to\mathcal{L}(-B)\to\mathcal{L}\to\mathcal{L}|_{B}\to 0,$
where $\mathcal{L}(-B):=\mathcal{L}\otimes\mathcal{O}_{X}(-B)$.
Since $\psi=\pi\circ f$ is projective, $f$ is projective. By Lemma 3.1 and
$\mathcal{L}(-B)-(\omega_{X}+\Delta-B)=\mathcal{L}-(\omega_{X}+\Delta),$
we have $R^{i}f_{*}(\mathcal{L}(-B))=0$ for every $i>0$.
Thus,
$0\to f_{*}\mathcal{L}(-B)\to f_{*}\mathcal{L}\to f_{*}(\mathcal{L}|_{B})\to 0$
is exact, and we obtain the long exact sequence
$\cdots\to R^{p}\pi_{*}(f_{*}\mathcal{L}(-B))\to R^{p}\pi_{*}(f_{*}\mathcal{L})\to R^{p}\pi_{*}(f_{*}(\mathcal{L}|_{B}))\to\cdots.$
By the definition of $B$, $f_{*}(\mathcal{L}|_{B})$ is either zero or a skyscraper sheaf. Therefore,
$R^{p}\pi_{*}(f_{*}(\mathcal{L}|_{B}))=0$
for every $p>0$. It remains to show that
$R^{p}\pi_{*}(f_{*}\mathcal{L}(-B))=0$
for every $p>0$.
Since $R^{i}f_{*}(\mathcal{L}(-B))=0$ for every $i>0$, the Leray spectral sequence yields
$R^{p}\pi_{*}(f_{*}\mathcal{L}(-B))=R^{p}\psi_{*}\mathcal{L}(-B)$
for every $p$.
Applying the vanishing theorem of Reid–Fukuda type (Lemma 3.1) to $\psi$, we obtain $R^{p}\psi_{*}\mathcal{L}(-B)=0$ for every $p>0$. Hence, $R^{p}\pi_{*}(f_{*}\mathcal{L}(-B))=0$ for every $p>0$, as desired. This completes the proof of Theorem 1.1. ∎
5. On the minimal model theory for log surfaces
In this section, we explain, paper by paper, that the minimal model theory for log surfaces developed in [F1], [F2], [F5], and [M] does not require the intricate vanishing theorems traditionally used in the literature. Theorem 1.1 and Lemma 3.2 already provide all the ingredients necessary for those arguments.
5.1. On the fundamental theorems established in [F1]
The minimal model theory discussed in [F1] depends on the torsion-free and vanishing theorem (see [F1, Theorem 6.3]). Let $f\colon Y\to X$ be a projective birational morphism from a smooth algebraic surface $Y$. In [F1, Theorem 6.3 (i)], it is shown that the torsion-free property, which is now called the strict support condition, holds for $f$. This is contained in Lemma 3.2 in the current paper. Moreover, [F1, Theorem 6.3 (ii)] easily follows from Theorem 1.1. In [F1, Section 8], the torsion-free and vanishing theorems are applied only in the case where $f\colon Y\to X$ is birational. Hence, the results in [F1, Section 8] for surfaces can be established solely by using Theorem 1.1 and Lemma 3.2. Thus, we recover [F1, Theorem 9.1] for surfaces.
We note the following well-known elementary property.
Lemma 5.1.
Let $(X,\Delta)$ be a log canonical surface and let $C$ be a minimal log canonical center of $(X,\Delta)$. If $\dim C=1$, then $(X,\Delta)$ is purely log terminal in a neighborhood of $C$. In particular, $C$ is normal and $(C,\Delta_{C})$ is kawamata log terminal, where $K_{C}+\Delta_{C}:=(K_{X}+\Delta)|_{C}$ by adjunction.
We provide a proof of Lemma 5.1 for completeness.
Proof of Lemma 5.1.
Assume that $\dim C=1$. Since $C$ is a minimal log canonical center by assumption, every divisor $E$ over $X$ whose center on $X$ intersects $C$ and is different from $C$ itself has discrepancy $a(E,X,\Delta)>-1$ by [F1, Theorem 9.1 (2)]. In particular, there is no divisor with discrepancy 1-1 whose center meets $C$ other than the divisor $C$. Hence $(X,\Delta)$ is purely log terminal in a neighborhood of $C$.
The remaining assertions follow immediately: since $(X,\Delta)$ is purely log terminal near $C$, by adjunction, the pair $(C,\Delta_{C})$ is kawamata log terminal. This completes the proof. ∎
By Lemma 5.1, when $X$ is a surface, if $W$ is a minimal log canonical center, then $(W,B_{W})$ is a kawamata log terminal curve, where $K_{W}+B_{W}:=(K_{X}+B)|_{W}$ by adjunction, or else $W$ is a point. Thus, [F1, Theorem 11.1] is clear when $X$ is a surface. In the proof of the non-vanishing theorem in [F1, Section 12], minimal log canonical centers that are disjoint from the non-lc locus are treated. By Lemma 5.1, such a center is either a point or a kawamata log terminal curve when the ambient variety is a surface. Therefore, no complicated arguments are required when working with surfaces. Thus, we can verify that Theorem 1.1 and Lemma 3.2 are sufficient to establish the results of [F1] for surfaces. We strongly recommend that the reader consult [F1, Section 2] for the appropriate use of the vanishing theorem for surfaces.
5.2. On the minimal model theory for log surfaces in [F2]
The Kodaira-type vanishing theorem stated in [F2, 2.8] follows immediately from Theorem 1.1. Note that the vanishing theorem in [F2, 2.8] is used repeatedly in [F2, Section 4]. In [F2, Section 9], which is an appendix, quasi-log structures on reducible curves are treated. We do not check here whether Theorem 1.1 and Lemma 3.2 are sufficient for the proof of [F2, Theorem 9.1]. We leave this as an exercise for the interested reader. Note that [F2, Theorem 9.1] is not necessary for the minimal model program and the abundance theorem for log surfaces (see [F2, Theorem 1.1]).
5.3. On the cone and contraction theorem established in [F5]
In [F5], the cone and contraction theorem is established for projective morphisms between complex analytic spaces. Note that [F5] is a complex analytic generalization of [F1]. The vanishing theorem (see [F5, Theorem 3.1.5]), which is one of the main results of [F6], can be recovered quickly from Theorem 1.1 and Lemma 3.2 when $f\colon X\to Y$ is a bimeromorphic morphism from a smooth complex analytic surface $X$. Therefore, [F5, Theorem 3.3.1] can be proved without difficulties when $X$ is a surface.
For the proof of the non-vanishing theorem (see [F5, Theorem 4.1.1]), no complicated vanishing theorems are needed if we treat only surfaces. In particular, we do not need [F5, Lemma 4.1.3]. This is because any minimal log canonical center of a log canonical surface is either a point or a kawamata log terminal curve by Lemma 5.1. Hence, in order to prove the results of [F5] for surfaces, Theorem 1.1 and Lemma 3.2 are sufficient.
5.4. On the minimal model theory for analytic log surfaces in [M]
In [M], we utilize the vanishing theorem stated in [M, Lemma 2.14]. It is not difficult to verify that the case where $f\colon\widetilde{X}\to X$ is a bimeromorphic morphism from a smooth complex analytic surface $\widetilde{X}$ is sufficient. In this case, [M, Lemma 2.14] can be recovered from Theorem 1.1. Thus, we can avoid the more intricate vanishing results proved in [F6] and [FF].
Acknowledgments.
The first author was partially supported by JSPS KAKENHI Grant Numbers JP21H04994 and JP23K20787. The authors are grateful to Yohei Hada for carefully reading the manuscript and kindly pointing out an error in an earlier version of this paper.